OpenStudy (anonymous):

Which of the following would increase the width of a confidence interval? Changing from a 99% to 95% confidence level. Increasing the variability of the outcome. Increasing the sample size. Removing an outlier from the data

OpenStudy (kropot72):

The confidence interval for the mean can be written \[\bar{x}-z\frac{\sigma}{\sqrt{n}}< \mu < \bar{x}+z\frac{\sigma}{\sqrt{n}}\] where the value of z is 2.576 to find a 99% confidence level and 1.960 to find a 95% confidence level. It can be seen that changing from a 99% to 95% confidence level would reduce the width of a confidence interval. What would happen to the width of a confidence interval if the variability of the outcome increased thus increasing the standard deviation?

OpenStudy (anonymous):

would increased. Thank you kropot

OpenStudy (kropot72):

You're welcome :)